Refer the table below. Our original information was that the probability of High demand was .6. We collected the following new information about what happened to demand when there was an improving or not improving economy. Answer the below question.

What is the probability of a not improving economy?

Improving Economy Not improving economy
Given high demand 0.77
Given Low demand 0.6

please, show the process
Answer #1

Probability of an improving economy given low demand = 1- Probability of a not improving economy given low demand

=1-0.6

=0.4

Probability of a not improving economy given high demand = 1- Probability of an improving economy given high demand

=1-0.77

=0.23

Hence probability of a not improving economy = 0.23*0.6+ 0.6*0.4

=0.378